# Electric Dipole Moments in 5+3 Flavor Weak Effective Theory

### v1

@@ 6.5 SMEFT Interpretation
The Yukawa couplings in SMEFT are more constrained as compared to the HEFT. Since, the leading non-standard effects to the Yukawa couplings are dim-6 (117), we can not have the double insertion of the Yukawas at this level. Hence, the contributions to the EDMs through flavor-violating Higgs couplings in Tab. 11 is essentially a dim-8 effect in the SMEFT power counting, because for the matching of $\big{[}{C}_{ff}^{S1,RR}\big{]}_{iijj}$ a double insertion of flavor-violating couplings are required. On the other hand, the EDMs through flavor-conserving couplings can be generated at dim-6 level in SMEFT. In this case, one of the Yukawa has to SM like. Moreover, since the SM Yukawas are restricted to be real in the mass basis, one can directly constrain the individual CP violating NP couplings from EDM observables unlike the generic case presented in Tab. 12.
@@
Limiting to dim-6 effects in the SMEFT, the matching onto 4f WET operators can be obtained by setting the one of the Yukawa to the SM values in (127)
$y^{d}_{ii,\textrm{SM}}={m_{d_{i}}\over v}\,,\quad y^{u}_{ii,\textrm{SM}}={m_{u_{i}}\over v}\,,\quad y^{e}_{ii,\textrm{SM}}={y_{e_{i}}\over v}.$ (147)
@@
Moreover, in the BZ contributions (6.1) we also set $\kappa_{V}=1$. The current limits on the CPV FC Higgs couplings are given in Tab. 14. For each case, we present two sets of results using the tree-level (4f) and two-loop (BZ) matching.
CPV FC Higgs couplings in SMEFT
current: $d_{n}$, Hg, Xe, Ra
matching Im($y^{d}_{11}$) Im($y^{d}_{22}$) Im($y^{d}_{33}$) Im($y^{u}_{11}$) Im($y^{u}_{22}$)
4f $(4.3\pm 2.3)\cdot 10^{-5}$ $(-5.0\pm 4.0)\cdot 10^{-4}$ $(-0.7\pm 2.4)\cdot 10^{-2}$ $(-2.7\pm 1.5)\cdot 10^{-5}$ $(0.0\pm 8.0)\cdot 10^{-2}$
BZ $(0.1\pm 1.5)\cdot 10^{-7}$ $(-1.0\pm 4.0)\cdot 10^{-4}$ $(0.0\pm 7.0)\cdot 10^{-3}$ $(-0.5\pm 3.4)\cdot 10^{-7}$ $(0.0\pm 5.0)\cdot 10^{-4}$
current: $\omega_{\rm HfF}$, $\omega_{\rm YbF}$, $\omega_{\rm ThO}$
matching Im($y^{d}_{11}$) Im($y^{d}_{22}$) Im($y^{d}_{33}$) Im($y^{u}_{11}$) Im($y^{u}_{22}$)
4f $(-2.0\pm 7.0)\cdot 10^{-7}$ $(-0.3\pm 1.0)\cdot 10^{-5}$ $(0.8\pm 1.3)\cdot 10^{-3}$ $(-2.0\pm 7.0)\cdot 10^{-7}$ $(3.0\pm 5.0)\cdot 10^{-4}$
@@
Table 14: The current limits on the CPV FC Yukawa couplings resulting due to tree-level and two-loop Barr-Zee diagrams up to dim-6 level in SMEFT. The neutron, Hg, Xe, and Ra EDMs and the frequencies $\omega_{\rm HfF}$, $\omega_{\rm YbF}$, $\omega_{\rm ThO}$ are included in the fit. The real parts of the Yukawas are set to their SM values.
These indirect bounds on the Higgs couplings are complementary to the constraints resulting from the direct Higgs searches as presented in the previous section.
@@ 7 Summary and Outlook
The inability of the SM to explain the origin of the baryon asymmetry in the Universe provides a strong motivation to search for new sources of CP violation beyond the phase of the CKM quark mixing matrix. Electric dipole moments of leptons, nucleons, atoms and molecules are extremely sensitive probes of flavor-diagonal CP violation. Upcoming experiments have the potential to improve the bounds on the neutron EDM and on EDMs of diamagnetic atoms by at least one order of magnitude, while molecular experiments will further constrain the electron EDM and start to probe hadronic EDMs. As they could offer the first hints of BSM physics, it is important to develop robust tools for the model-independent interpretation of EDM experiments. In addition, identifying the detailed features of BSM physics from one or multiple observations in EDM experiments requires understanding their correlations with observables in other systems and at other energy scales. Effective Field Theories provide a natural framework to achieve these goals in a controlled and systematic way, with minimal reliance on specific BSM models. In this paper we discuss EDMs in the Weak EFT, and leverage EFT tools such as renormalization group evolution and matching at the heavy fermion thresholds to derive master formulae for EDMs in terms of the complete set of $\Delta F=0$ WET operators at the electroweak scale. This set includes both operators with light $u$, $d$ and $s$ quarks, which have nonvanishing nucleon matrix elements and thus contribute to EDM at “tree level”, but also operators with heavy $b$ and $c$ quarks and leptons of the second and third generations. Indeed we find that the great majority of CP-violating $\Delta F=0$ operators with heavy fermions mixes into “tree level” operators at leading or next to leading log. EDM constraints are so strong that even in the case of operators that contribute at next-to-leading log (meaning two or even more loops) the operator scale has to be larger than 1 TeV. We now briefly summarize the main results of this paper.
@@
  • The main result of this paper are the master formulae for the HfF, ThO and YbF precession frequencies, given in Tables 7 and 15, and for the neutron, proton and Hg EDMs, in Tables 8, 9, 10 and 16. These formulae can be readily applied to EFTs at the electroweak scale (SMEFT or HEFT) or to BSM models by calculating the WET matching coefficients at the electroweak scale.
@@
The coefficients $\alpha_{I}^{X}(\mu_{ew})$ provided in these tables encode both short-distance effects, from the running between the electroweak scale $\mu_{ew}$ and the low energy scales $\mu_{low}=2$ GeV (or the heavy fermion thresholds), and long distance effects, from hadronic and nuclear matrix elements. Concerning the short distance effects:
@@
  • We consider the full WET one-loop anomalous dimension, which is sufficient to capture the LL contributions from scalar four-quark operators with two heavy and two light quarks ( $\big{[}{C}_{uu}^{S1(8),RR}\big{]}$, $\big{[}{C}_{ud}^{S1(8),RR}\big{]}$, and $\big{[}{C}_{uddu}^{S1(8),RR}\big{]}$), scalar leptonic operators with two heavy and two light leptons ($\big{[}{C}_{ee}^{S,RR}\big{]}$) and scalar and tensor semileptonic operators with light quarks and heavy leptons or light leptons and heavy quarks ($\big{[}{C}_{eu(d)}^{T,RR}\big{]}$ and $\big{[}{C}_{eu(d)}^{S,RR}\big{]}$). The one-loop RGE and the one-loop matching of heavy quark QCD dipoles onto the Weinberg three-gluon operator provide NLL contributions from four-quark scalar operators with four heavy quarks, and semileptonic scalar and tensor operators with heavy quarks and heavy leptons.
@@
  • The vector operators $\big{[}{C}_{uu}^{V1(8),LR}\big{]}$, $\big{[}{C}_{dd}^{V1(8),LR}\big{]}$, $\big{[}{C}_{uddu}^{V1(8),LR}\big{]}$ and $\big{[}{C}_{ee}^{V,LR}\big{]}$, with two or four heavy fermions, do not generate LL contributions. For these operators, we identify NLL contributions by generalizing two-loop anomalous dimensions originally derived for $B\rightarrow X_{s}\gamma$ Cho:1993zb , and by considering the one loop matching at the heavy fermion threshold. After two loop running and one loop matching onto the $b$ and $c$ QCD dipole, $\big{[}{C}_{uddu}^{V1(8),LR}\big{]}_{2332}$ generates a NNLL contribution to EDMs.
@@
As an example, we have applied the master formulae to study the constraints from EDMs on Higgs couplings. For flavor-violating Higgs couplings, we showed that, at the moment, $\Delta F=2$ observables provide stronger bounds. Future EDM experiments, however, will provide competitive constraints, especially in the $bd$ sector. In the case of flavor conserving Yukawa couplings, we updated bounds on the imaginary part in two scenarios. In the HEFT scenario, in which the real part of the Yukawa is also a free parameter, we found that EDMs give the strongest bounds on the imaginary part of the $u$, $d$ and $c$ Yukawas, while Higgs observables dominate the bounds on the $s$ and $b$ Yukawas. In the SMEFT scenario, in which, up to dimension-eight corrections, the real part of the Yukawa is fixed to its SM value, EDMs give by far the strongest bounds. In particular, the electron EDM gives the strongest bounds on the $b$ and $s$ Yukawas, neutron and mercury dominate the $u$ and $d$ bounds, and for the charm quark both contributions are important.
@@
This work can be extended in several directions. First of all, we have in this paper relied on a logarithmic counting to organize the contributions to EDMs arising from the solution of the WET RGEs at one- and two-loops. This counting does not take into account the presence of small couplings, such as the QED coupling, for which the suppression from additional loops is not fully offset by the large logarithms. It is thus possible that higher loop corrections to matching and running, while formally subleading, will be numerically more important that the LL and NLL terms identified here, especially in those cases in which these effectively arise at three loops (e.g. $\big{[}{C}_{uddu}^{V1(8),LR}\big{]}_{2332}$ or $\big{[}{C}_{ud}^{S1(8),RR}\big{]}_{2233}$). Especially for applications to very sensitive observables as EDMs, it is thus important to extend the calculations of WET anomalous dimensions and threshold corrections to higher order.
@@
Secondly, higher order corrections are necessary to constrain the few $\Delta F=0$ CP-violating operators that are not listed in Table 1. These include $\big{[}{C}_{eu}^{S,RL}\big{]}_{1122}$, $\big{[}{C}_{eu}^{S,RL}\big{]}_{1133}$ and all the entries with heavy leptons, for which it should be possible to generalize the two-loop anomalous dimension for vector operators, and $\big{[}{C}_{ee}^{S,RR}\big{]}_{iijj}$ with $i,j>1$. In addition, we find very weak constraints on $\mu$-charm and $\tau$-charm semileptonic operators ($\big{[}{C}_{eu}^{T(S),RR}\big{]}_{2222}$, $\big{[}{C}_{eu}^{T(S),RR}\big{]}_{3322}$) and $\mu$-bottom and $\tau$-bottom semileptonic operators ($\big{[}{C}_{ed}^{T(S),RR}\big{]}_{2233}$ and $\big{[}{C}_{ed}^{T(S),RR}\big{]}_{3333}$). For these classes, it will be important to find additional paths.
@@
Finally, reducing the theoretical uncertainties is crucial to the success of the EDM program. This requires, in particular, to address the large hadronic and nuclear uncertainties that affect the master formulae provided in this paper.
@@ Acknowledgements
We acknowledge several useful discussions with W. Dekens and M. Misiak. JK and EM are supported by the U.S. Department of Energy through the Los Alamos National Laboratory and by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under project numbers 20220706PRD1 and 20240078DR. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001).
@@ Appendix A Additional Formulae
$[C_{e\gamma}]_{11}$ $(10.0\pm 2.1)\cdot 10^{12}$ $[C^{S,RR}_{eu}]_{1122}$ $(-32.2\pm 6.0)\cdot 10^{5}$
$[C^{S,RR}_{ee}]_{1111}$ $(9.4\pm 1.6)\cdot 10^{4}$ $[C^{T,RR}_{eu}]_{1122}$ $(13.3\pm 2.5)\cdot 10^{8}$
$[C^{S,RR}_{eu}]_{1111}$ $(14.6\pm 3.1)\cdot 10^{8}$ $[C^{S,RR}_{ed}]_{1133}$ $(-11.5\pm 2.3)\cdot 10^{5}$
$[C^{S,RR}_{ed}]_{1111}$ $(14.9\pm 2.9)\cdot 10^{8}$ $[C^{T,RR}_{ed}]_{1133}$ $(-16.1\pm 3.1)\cdot 10^{8}$
$[C^{S,RR}_{ed}]_{1122}$ $(11.0\pm 2.2)\cdot 10^{7}$ $[C^{S,RR}_{ee}]_{1122}$ $(-20.2\pm 4.4)\cdot 10^{4}$
$[C^{S,RL}_{eu}]_{1111}$ $(14.5\pm 3.2)\cdot 10^{8}$ $[C^{S,RR}_{ee}]_{1133}$ $(-44.1\pm 8.1)\cdot 10^{5}$
$[C^{S,RL}_{ed}]_{1111}$ $(14.9\pm 3.0)\cdot 10^{8}$ $[C^{S,RR}_{ee}]_{1221}$ $(9.9\pm 1.9)\cdot 10^{6}$
$[C^{S,RL}_{ed}]_{1122}$ $(10.9\pm 2.1)\cdot 10^{7}$ $[C^{S,RR}_{ee}]_{1331}$ $(19.1\pm 3.5)\cdot 10^{7}$
$[C^{T,RR}_{eu}]_{1111}$ $(-8.7\pm 1.6)\cdot 10^{8}$ $[C^{V,LR}_{ee}]_{1221}$ $(-25.2\pm 5.4)\cdot 10^{4}$
$[C^{T,RR}_{ed}]_{1111}$ $(43.8\pm 8.0)\cdot 10^{7}$ $[C^{V,LR}_{ee}]_{1331}$ $(-25.2\pm 4.6)\cdot 10^{4}$
$[C^{T,RR}_{ed}]_{1122}$ $(31.4\pm 6.6)\cdot 10^{7}$
@@
$[C_{e\gamma}]_{11}$ $(16.1\pm 1.2)\cdot 10^{11}$ $[C^{S,RR}_{eu}]_{1122}$ $(-52.2\pm 3.7)\cdot 10^{4}$
$[C^{S,RR}_{ee}]_{1111}$ $(15.2\pm 1.3)\cdot 10^{3}$ $[C^{T,RR}_{eu}]_{1122}$ $(21.6\pm 1.6)\cdot 10^{7}$
$[C^{S,RR}_{eu}]_{1111}$ $(14.1\pm 2.0)\cdot 10^{7}$ $[C^{S,RR}_{ed}]_{1133}$ $(-18.7\pm 1.4)\cdot 10^{4}$
$[C^{S,RR}_{ed}]_{1111}$ $(14.3\pm 2.1)\cdot 10^{7}$ $[C^{T,RR}_{ed}]_{1133}$ $(-26.1\pm 2.0)\cdot 10^{7}$
$[C^{S,RR}_{ed}]_{1122}$ $(10.7\pm 1.3)\cdot 10^{6}$ $[C^{S,RR}_{ee}]_{1122}$ $(-32.8\pm 2.7)\cdot 10^{3}$
$[C^{S,RL}_{eu}]_{1111}$ $(14.1\pm 2.0)\cdot 10^{7}$ $[C^{S,RR}_{ee}]_{1133}$ $(-71.4\pm 5.3)\cdot 10^{4}$
$[C^{S,RL}_{ed}]_{1111}$ $(14.3\pm 2.3)\cdot 10^{7}$ $[C^{S,RR}_{ee}]_{1221}$ $(16.0\pm 1.3)\cdot 10^{5}$
$[C^{S,RL}_{ed}]_{1122}$ $(10.5\pm 1.3)\cdot 10^{6}$ $[C^{S,RR}_{ee}]_{1331}$ $(31.0\pm 2.3)\cdot 10^{6}$
$[C^{T,RR}_{eu}]_{1111}$ $(-130.2\pm 8.9)\cdot 10^{6}$ $[C^{V,LR}_{ee}]_{1221}$ $(-40.8\pm 3.3)\cdot 10^{3}$
$[C^{T,RR}_{ed}]_{1111}$ $(65.2\pm 4.5)\cdot 10^{6}$ $[C^{V,LR}_{ee}]_{1331}$ $(-40.8\pm 3.2)\cdot 10^{3}$
$[C^{T,RR}_{ed}]_{1122}$ $(50.5\pm 4.1)\cdot 10^{6}$

### v2

@@ 6.5 SMEFT Interpretation
The Yukawa couplings in SMEFT are more constrained as compared to the HEFT. Since the leading non-standard effects to the Yukawa couplings are dim-6 (117), we can not have the double insertion of the Yukawas at this level. Hence, the contributions to the EDMs through flavor-violating Higgs couplings in Tab. 11 is essentially a dim-8 effect in the SMEFT power counting, because for the matching of $\big{[}{C}_{ff}^{S1,RR}\big{]}_{iijj}$ a double insertion of flavor-violating couplings is required. On the other hand, flavor-conserving CPV Yukawa couplings generate EDMs at dim-6 in SMEFT. In this case, one of the Yukawa has to be SM-like. Moreover, since the SM Yukawas are restricted to be real in the mass basis, one can directly constrain the individual CPV couplings from EDM observables unlike the generic case presented in Tab. 12.
@@
Limiting to dim-6 effects in the SMEFT, the matching onto 4f WET operators can be obtained by setting one of the Yukawas in (127) to its SM value
$y^{d}_{ii,\textrm{SM}}={m_{d_{i}}\over v}\,,\quad y^{u}_{ii,\textrm{SM}}={m_{u_{i}}\over v}\,,\quad y^{e}_{ii,\textrm{SM}}={m_{e_{i}}\over v}.$ (145)
@@
Moreover, in the BZ contributions (6.1) we also set $\kappa_{V}=1$. The current limits on the CPV FC Higgs couplings are given in Tab. 14. For each case, we present two sets of results using the tree-level (4f) and 2-loop (BZ) matching. Also, in this case, we neglect theoretical errors. Once they are under better control, theoretical errors can straightforwardly be added to the analysis.
CPV FC Higgs couplings in SMEFT at $1\sigma$ level
$d_{n}$, Hg, Xe, Ra
[…]
4f $1.4\cdot 10^{-5}$ $2.6\cdot 10^{-3}$ $2.6\cdot 10^{-3}$ $6.0\cdot 10^{-6}$ $4.0\cdot 10^{-3}$
BZ $9.0\cdot 10^{-8}$ $7.0\cdot 10^{-4}$ $9.0\cdot 10^{-3}$ $1.1\cdot 10^{-7}$ $1.3\cdot 10^{-3}$
$\omega_{\rm HfF}$, $\omega_{\rm YbF}$, $\omega_{\rm ThO}$
[…]
4f $4.0\cdot 10^{-3}$ $4.0\cdot 10^{-3}$ $2.0\cdot 10^{-3}$ $9.0\cdot 10^{-4}$ $8.0\cdot 10^{-4}$ $2.6\cdot 10^{-8}$ $1.6\cdot 10^{-3}$ $1.0\cdot 10^{-4}$
BZ $1.5\cdot 10^{-9}$
combined
[…]
4f +BZ $9.0\cdot 10^{-8}$ $7.0\cdot 10^{-4}$ $1.8\cdot 10^{-3}$ $1.0\cdot 10^{-7}$ $7.0\cdot 10^{-4}$ $1.5\cdot 10^{-9}$ $1.6\cdot 10^{-3}$ $1.0\cdot 10^{-4}$
@@
Table 14: The current limits on the CPV FC Yukawa couplings resulting due to tree-level and 2-loop Barr-Zee diagrams up to dim-6 level in SMEFT. The neutron, Hg, Xe, and Ra EDMs and the frequencies $\omega_{\rm HfF}$, $\omega_{\rm YbF}$, $\omega_{\rm ThO}$ are included in the fit. The real parts of the Yukawas are set to their SM values. The matching scale is set to the Higgs mass.
The imaginary part of the Yukawa couplings of the $u$ and $d$ quarks is predominantly constrained by the neutron and Hg EDMs, via the contributions arising from Barr-Zee diagrams. The constraints are very strong, and out of the reach of existing colliders. The $s$ quark is also mostly constrained by hadronic EDMs. In this case, the bound is a factor of ten stronger than the constraints from Higgs observables discussed in Sec. 6.4, so that, once theoretical errors are considered, the LHC and EDM experiments can probe similar scales. The $b$-quark Yukawa is mostly constrained by the electron EDM. These contributions arise from first integrating out the Higgs and matching onto the scalar four-fermion operator $\big{[}{C}_{ed}^{S,RR}\big{]}$, which consequently runs into $\big{[}{C}_{ed}^{T,RR}\big{]}$ and $C_{e\gamma}$. Finally, the $c$-quark Yukawa receive similar constraints from $d_{n}$, $d_{\rm Hg}$ and $\omega_{\rm HfF}$. For the $c$ and $b$ Yukawas, the indirect bounds on the Higgs couplings are complementary to the constraints resulting from the direct Higgs searches as presented in the previous section.
The bounds in Table 14 are of the same order of magnitude as those presented in Ref. Chien:2015xha . The main difference is in the treatment of the RGE between the EW and $\sim$ 2 GeV scale. Ref. Chien:2015xha ignored the role of four-fermion operators, which re-appeared as unresummed large logarithms of $m_{q}/m_{h}$ in the matching coefficients. Our analysis is more similar to Ref. Brod:2023wsh , which, for the $b$ and $c$ quarks, calculated the matching coefficients and the mixing between $C_{e\gamma}$, $\big{[}{C}_{eq}^{S,RR}\big{]}$ and $\big{[}{C}_{eq}^{T,RR}\big{]}$ at one order higher than in our study. The bounds on $y^{d}_{33}$ and $y^{u}_{22}$ obtained in Ref. Brod:2023wsh are very close to those presented in Table 14.
@@ 7 Summary and Outlook
The inability of the SM to explain the origin of the baryon asymmetry in the Universe provides a strong motivation to search for new sources of CP violation beyond the phase of the CKM quark mixing matrix. Electric dipole moments of leptons, nucleons, atoms, and molecules are extremely sensitive probes of flavor-diagonal CP violation. Upcoming experiments have the potential to improve the bounds on the neutron EDM and on EDMs of diamagnetic atoms by at least one order of magnitude, while molecular experiments will further constrain the electron EDM and start to probe hadronic EDMs. As they could offer the first hints of BSM physics, it is important to develop robust tools for the model-independent interpretation of EDM experiments. In addition, identifying the detailed features of BSM physics from one or multiple observations in EDM experiments requires understanding their correlations with observables in other systems and at other energy scales. Effective Field Theories provide a natural framework to achieve these goals in a controlled and systematic way, with minimal reliance on specific BSM models. In this paper, we discuss EDMs in the Weak EFT and leverage EFT tools such as renormalization group evolution and matching at the heavy fermion thresholds to derive master formulae for EDMs in terms of the complete set of $\Delta F=0$ WET operators at the electroweak scale. This set includes both operators with light $u$, $d$, and $s$ quarks, which have nonvanishing nucleon matrix elements and thus contribute to EDM at “tree level”, but also operators with heavy $b$ and $c$ quarks and leptons of the second and third generations. Indeed we find that the great majority of CP-violating $\Delta F=0$ operators with heavy fermions mix into tree-level operators at leading or next to leading-log. EDM constraints are so strong that even in the case of operators that contribute at next-to-leading log (meaning two or even more loops) the operator scale has to be larger than 1 TeV. We now briefly summarize the main results of this paper.
@@
  • The main result of this paper is the master formulae for the HfF, ThO, and YbF precession frequencies, given in Tables 7 and 15, and for the neutron, proton, and Hg EDMs, in Tables 8, 9, 10 and 16. These formulae can be readily applied to EFTs at the electroweak scale (SMEFT or HEFT) or to BSM models by calculating the WET matching coefficients at the electroweak scale.
@@
The coefficients $\alpha_{I}^{X}(\mu_{ew})$ provided in these tables encode both short-distance effects, from the running between the electroweak scale $\mu_{ew}$ and the low energy scales $\mu_{low}=2$ GeV (or the heavy fermion thresholds), and long-distance effects, from hadronic and nuclear matrix elements. Concerning the short-distance effects:
@@
  • We consider the full WET 1-loop anomalous dimension, which is sufficient to capture the LL contributions from scalar four-quark operators with two heavy and two light quarks ( $\big{[}{C}_{uu}^{S1(8),RR}\big{]}$, $\big{[}{C}_{ud}^{S1(8),RR}\big{]}$, and $\big{[}{C}_{uddu}^{S1(8),RR}\big{]}$), scalar leptonic operators with two heavy and two light leptons ($\big{[}{C}_{ee}^{S,RR}\big{]}$) and scalar and tensor semileptonic operators with light quarks and heavy leptons or light leptons and heavy quarks ($\big{[}{C}_{eu(d)}^{T,RR}\big{]}$ and $\big{[}{C}_{eu(d)}^{S,RR}\big{]}$). The 1-loop RGE and the 1-loop matching of heavy quark QCD dipoles onto the Weinberg three-gluon operator provide NLL contributions from four-quark scalar operators with four heavy quarks, and semileptonic scalar and tensor operators with heavy quarks and heavy leptons.
@@
  • The vector operators $\big{[}{C}_{uu}^{V1(8),LR}\big{]}$, $\big{[}{C}_{dd}^{V1(8),LR}\big{]}$, $\big{[}{C}_{uddu}^{V1(8),LR}\big{]}$ and $\big{[}{C}_{ee}^{V,LR}\big{]}$, with two or four heavy fermions, do not generate LL contributions. For these operators, we identify NLL contributions by generalizing 2-loop anomalous dimensions originally derived for $B\rightarrow X_{s}\gamma$ Cho:1993zb , and by considering the 1-loop matching at the heavy fermion threshold. After 2-loop running and 1-loop matching onto the $b$ and $c$ QCD dipole, $\big{[}{C}_{uddu}^{V1(8),LR}\big{]}_{2332}$ generates a NNLL contribution to EDMs.
@@
As an example, we have applied the master formulae to study the constraints from EDMs on Higgs couplings. For flavor-violating Higgs couplings, we showed that, at the moment, $\Delta F=2$ observables provide stronger bounds. Future EDM experiments, however, will provide competitive constraints, especially in the $bd$ sector. In the case of flavor conserving Yukawa couplings, we updated bounds on the imaginary part in two scenarios. In the HEFT scenario, in which the real part of the Yukawa is also a free parameter, we found that EDMs give the strongest bounds on the imaginary part of the $e$, $u$, and $d$ Yukawas, while Higgs observables dominate the bounds on the $s$, $c$ and $b$ Yukawas. In the SMEFT scenario, in which, up to dimension-eight corrections, the real part of the Yukawa is fixed to its SM value, EDMs give by far the strongest bounds. In particular, the electron EDM gives the strongest bounds on the $b$ and $s$ Yukawas, neutron and mercury dominate the $u$ and $d$ bounds, and for the charm quark both contributions are important.
@@
This work can be extended in several directions. First of all, we have in this paper relied on a logarithmic counting to organize the contributions to EDMs arising from the solution of the WET RGEs at one and two loops. This counting does not take into account the presence of small couplings, such as the QED coupling, for which the suppression from additional loops is not fully offset by the large logarithms. It is thus possible that higher loop corrections to matching and running, while formally subleading, will be numerically more important that the LL and NLL terms identified here, especially in those cases in which these effectively arise at three loops (e.g. $\big{[}{C}_{uddu}^{V1(8),LR}\big{]}_{2332}$ or $\big{[}{C}_{ud}^{S1(8),RR}\big{]}_{2233}$). Especially for applications to very sensitive observables as EDMs, it is thus important to extend the calculations of WET anomalous dimensions and threshold corrections to higher order.
@@
Secondly, higher order corrections are necessary to constrain the few $\Delta F=0$ CP-violating operators that are not listed in Table 1. These include $\big{[}{C}_{eu}^{S,RL}\big{]}_{1122}$, $\big{[}{C}_{eu}^{S,RL}\big{]}_{1133}$ and all the entries with heavy leptons, for which it should be possible to generalize the 2-loop anomalous dimension for vector operators, and $\big{[}{C}_{ee}^{S,RR}\big{]}_{iijj}$ with $i,j>1$. In addition, we find very weak constraints on $\mu$-charm and $\tau$-charm semileptonic operators ($\big{[}{C}_{eu}^{T(S),RR}\big{]}_{2222}$, $\big{[}{C}_{eu}^{T(S),RR}\big{]}_{3322}$) and $\mu$-bottom and $\tau$-bottom semileptonic operators ($\big{[}{C}_{ed}^{T(S),RR}\big{]}_{2233}$ and $\big{[}{C}_{ed}^{T(S),RR}\big{]}_{3333}$). For these classes, it will be important to find additional paths.
@@
Finally, reducing the theoretical uncertainties is crucial to the success of the EDM program. This requires, in particular, to address the large hadronic and nuclear uncertainties that affect the master formulae provided in this paper. In future the theoretical uncertaintities can be included in the fits using smelli program Aebischer:2018iyb .
@@ Acknowledgements
We acknowledge several useful discussions with W. Dekens and M. Misiak. JK and EM are supported by the U.S. Department of Energy through the Los Alamos National Laboratory and by the Laboratory Directed Research and Development program of Los Alamos National Laboratory under project numbers 20220706PRD1 and 20240078DR. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). We acknowledge support from the DOE Topical Collaboration “Nuclear Theory for New Physics,” award No. DE-SC0023663.
@@ Appendix A Additional Formulae
$[C_{e\gamma}]_{11}$ $(9.9\pm 2.1)\cdot 10^{12}$ $[C^{S,RR}_{eu}]_{1122}$ $(-3.2\pm 0.6)\cdot 10^{6}$
$[C^{S,RR}_{ee}]_{1111}$ $(9.4\pm 2.1)\cdot 10^{4}$ $[C^{T,RR}_{eu}]_{1122}$ $(1.3\pm 0.3)\cdot 10^{9}$
$[C^{S,RR}_{eu}]_{1111}$ $(1.5\pm 0.3)\cdot 10^{9}$ $[C^{S,RR}_{ed}]_{1133}$ $(-1.3\pm 0.2)\cdot 10^{6}$
$[C^{S,RR}_{ed}]_{1111}$ $(1.5\pm 0.3)\cdot 10^{9}$ $[C^{T,RR}_{ed}]_{1133}$ $(-1.7\pm 0.3)\cdot 10^{9}$
$[C^{S,RR}_{ed}]_{1122}$ $(1.1\pm 0.2)\cdot 10^{8}$ $[C^{S,RR}_{ee}]_{1122}$ $(-2.0\pm 0.4)\cdot 10^{5}$
$[C^{S,RL}_{eu}]_{1111}$ $(1.5\pm 0.3)\cdot 10^{9}$ $[C^{S,RR}_{ee}]_{1133}$ $(-4.4\pm 0.7)\cdot 10^{6}$
$[C^{S,RL}_{ed}]_{1111}$ $(1.5\pm 0.3)\cdot 10^{9}$ $[C^{S,RR}_{ee}]_{1221}$ $(9.9\pm 2.0)\cdot 10^{6}$
$[C^{S,RL}_{ed}]_{1122}$ $(1.1\pm 0.3)\cdot 10^{8}$ $[C^{S,RR}_{ee}]_{1331}$ $(1.9\pm 0.4)\cdot 10^{8}$
$[C^{T,RR}_{eu}]_{1111}$ $(-8.7\pm 1.6)\cdot 10^{8}$ $[C^{V,LR}_{ee}]_{1221}$ $(-3.2\pm 0.6)\cdot 10^{6}$
$[C^{T,RR}_{ed}]_{1111}$ $(4.4\pm 0.8)\cdot 10^{8}$ $[C^{V,LR}_{ee}]_{1331}$ $(-5.1\pm 1.0)\cdot 10^{7}$
$[C^{T,RR}_{ed}]_{1122}$ $(3.1\pm 0.5)\cdot 10^{8}$
@@
$[C_{e\gamma}]_{11}$ $(1.6\pm 0.1)\cdot 10^{12}$ $[C^{S,RR}_{eu}]_{1122}$ $(-5.2\pm 0.4)\cdot 10^{5}$
$[C^{S,RR}_{ee}]_{1111}$ $(1.5\pm 0.1)\cdot 10^{4}$ $[C^{T,RR}_{eu}]_{1122}$ $(2.1\pm 0.2)\cdot 10^{8}$
$[C^{S,RR}_{eu}]_{1111}$ $(1.4\pm 0.2)\cdot 10^{8}$ $[C^{S,RR}_{ed}]_{1133}$ $(-2.1\pm 0.2)\cdot 10^{5}$
$[C^{S,RR}_{ed}]_{1111}$ $(1.4\pm 0.2)\cdot 10^{8}$ $[C^{T,RR}_{ed}]_{1133}$ $(-2.7\pm 0.2)\cdot 10^{8}$
$[C^{S,RR}_{ed}]_{1122}$ $(1.1\pm 0.1)\cdot 10^{7}$ $[C^{S,RR}_{ee}]_{1122}$ $(-3.3\pm 0.2)\cdot 10^{4}$
$[C^{S,RL}_{eu}]_{1111}$ $(1.4\pm 0.2)\cdot 10^{8}$ $[C^{S,RR}_{ee}]_{1133}$ $(-7.1\pm 0.5)\cdot 10^{5}$
$[C^{S,RL}_{ed}]_{1111}$ $(1.4\pm 0.2)\cdot 10^{8}$ $[C^{S,RR}_{ee}]_{1221}$ $(1.6\pm 0.1)\cdot 10^{6}$
$[C^{S,RL}_{ed}]_{1122}$ $(1.1\pm 0.1)\cdot 10^{7}$ $[C^{S,RR}_{ee}]_{1331}$ $(3.1\pm 0.2)\cdot 10^{7}$
$[C^{T,RR}_{eu}]_{1111}$ $(-1.3\pm 0.1)\cdot 10^{8}$ $[C^{V,LR}_{ee}]_{1221}$ $(-5.1\pm 0.4)\cdot 10^{5}$
$[C^{T,RR}_{ed}]_{1111}$ $(6.5\pm 0.5)\cdot 10^{7}$ $[C^{V,LR}_{ee}]_{1331}$ $(-8.3\pm 0.7)\cdot 10^{6}$
$[C^{T,RR}_{ed}]_{1122}$ $(5.1\pm 0.4)\cdot 10^{7}$